A particular cylindrical bucket has a height of 50.0 cm, and the radius of its circular cross-section is 15 cm. The bucket is empty, aside from containing air. The bucket is then inverted so that its open end is down and, being careful not to lose any of the air trapped inside, the bucket is lowered below the surface of a fresh-water lake so the water-air interface in the bucket is 40.0 m below the surface of the lake. To keep the calculations simple, use g = 10 m/s^2, atmospheric pressure = 1 x 10^5 Pa, and the density of water as 1000 kg/m^3.
If the temperature is the same at the surface of the lake and at a depth of 40.0 m below the surface , what is the height of the cylinder of air in the bucket when the bucket is at a depth of 40.0 m below the surface of the lake?
_______ cm

A particular cylindrical bucket has a height of 50.0 cm, and the radius of its circular cross-section is 15 cm. The bucket is empty, aside from containing air. The bucket is then inverted so that its open end is down and, being careful not to lose any of the air trapped inside, the bucket is lowered below the surface of a fresh-water lake so the water-air interface in the bucket is 40.0 m below the surface of the lake. To keep the calculations simple, use g = 10 m/s^2, atmospheric pressure = 1 x 10^5 Pa, and the density of water as 1000 kg/m^3.
If, instead, the temperature changes from 300K at the surface of the lake to 275K at a depth of 40.0 m below the surface of the lake, what is the height of the cylinder of air in the bucket?
_______ cm
Answer #1

Temperature is same at surface and also at 40 m depth

we have Pi Vi = Pf Vf ........(1)

where Pi = initial pressure at surface of lake = 1\times105 Pa

Vi = initial volume = \pi \times152\times 50 cm3

Pf = pressure at depth of 40 m = 1\times105 + \rho gh = 1\times105 + 1000\times10\times40 = 5 \times 105 Pa

Hence from eqn.(1), we get Vf = ( Pi / Pf ) Vi = (1/5) \times \pi\times152\times 50 = \pi \times152\times 10 cm3

Water height inside bucket = 10 cm

-------------------------------------------------------------------------------------

Temperature at surface of lake = 300 K ; Temperature at 40 m depth = 275 K

we have, ( Pi Vi ) / Ti = ( Pf Vf ) / Tf   .....................(2)

Vf = ( Pi / Pf ) \times ( Tf / Ti ) \times Vi = (1/5) \times (275/300) \times \pi\times152\times 50

height of water in the bucket, when it is at 40 m deep inside lake = (1/5) \times (275/300) \times 50 = 9.17 cm